OpenStudy (anonymous):

[(3-i)^1/2]/1+i=x+yi... i for imaginery number, help me pls, my brain is bursting...

OpenStudy (anonymous):

The question is...?

OpenStudy (anonymous):

just solve this equation, find the value of x and y as they are real number

OpenStudy (anonymous):

Take the square of both sides, the left hand side will be: \[{3-i \over (1+i)^2}={3-i \over 2i}=-{1 \over 2}-{3 \over 2}i\] The right hand side will be \((x+yi)^2=x^2+ x y i-y^2\). So, w have the two equations: \(x^2-y^2=-\frac{1}{2}\) and \(xy=-\frac{3}{2}\). Solve them and get your values.

OpenStudy (anonymous):

Don't forget to check each value you get and then reject all values that don't satisfy the original equation.

OpenStudy (anonymous):

It should be \(2xy=-3/2\)